Write an algebraic equation for the english statement. "Subtract 5 from the sum of x and 7 times 3 and the result is 20.
I came up with -5+6 = 7x3=20
I don't know if this is written in the right equation, please help!

Answer by praseenakos@yahoo.com(507)

Question:

Write an algebraic equation for the english statement. "Subtract 5 from the sum of x and 7 times 3 and the result is 20.
I came up with -5+6 = 7x3=20

Answer:

sum of x and 7 = x + 7

sum of x and 7 times 3 = (x + 7 ) * 3

Subtract 5 from the sum of x and 7 times 3 = (x + 7)3 - 5

Subtract 5 from the sum of x and 7 times 3 and the result is 20 ==>

==> (x + 7)3 - 5 = 20


or you can write, 3 (x + 7) - 5 = 20

hope you found the explanation useful.

Regards.

Praseena.